# Predict-then-Optimize Method for Seaport Power–Logistics Scheduling: Generalization across Varying Tasks Stream

### v1

@@ Abstract
Power–logistics scheduling in modern seaports typically follow a predict-then-optimize pipeline. To enhance decision quality, decision-focused learning has been proposed to align forecasting and optimization via end-to-end training. However, most formulations assume a fixed task configuration in downstream optimization, and thus generalize poorly to evolving task structures induced by varying seaport vessel arrivals. We address this gap with a decision-focused continual learning framework that adapts online to a stream of scheduling tasks. Specifically, we introduce Fisher information based regularization to enhance cross-task generalization by preserving parameters critical to prior tasks. A differentiable convex surrogate is also developed to stabilize gradient backpropagation. The proposed approach enables learning a decision-aligned forecasting model for new scheduling tasks while retaining generalization on earlier tasks. Experiments calibrated to the Jurong Port demonstrate superior decision performance and generalization over existing methods with reduced computational cost.
@@
[2]organization=School of Electrical and Electronic Engineering, Nanyang Technological University, city=Singapore, country=Singapore
@@ 1 Introduction
To address this misalignment, decision-focused learning (DFL) instead evaluates the decision quality of forecasts through their impact on the specific downstream optimization tasks [8]. As depicted in Fig. 1(b), DFL integrates forecasting and optimization in an end-to-end pipeline and trains the forecasting model directly by the decision-related loss. Recent work further refines DFL for power system applications by deriving task-aware losses and differentiable solvers that facilitate stable end-to-end training. These studies have established an effective learning paradigm for enhancing decision-making benefits in various power system contexts, including unit commitment [9, 10], economic dispatch [12, 13, 14], cost-oriented renewable-energy forecasting and trading [15, 16, 17], energy-storage arbitrage [18, 19], and multi-energy scheduling [20].
However, existing DFL settings focus on learning a forecasting model dedicated to optimization tasks with fixed and predetermined structures. In contrast, PLS scheduling tasks are continuously evolving due to varying vessel arrivals, since logistics decisions must be made conditional on vessel arrivals as exogenous parameters [7]. Each new arrival configuration introduces a distinct task instance characterized by different numbers of vessels, cargo profiles, and time-coupled constraints. As a result, although these tasks share similar high-level formulations, they differ in the dimensions of decision variables and the geometry of feasible regions. Consequently, the ground-truth objectives and loss landscapes of DFL is correspondingly dynamic across different configurations. This constitutes the fundamental distinction between PLS operation and conventional DFL settings.
@@
Online continual learning thus offers a promising direction, yet two key challenges still remain: i) How to retain decision effectiveness of the forecasting model on past tasks during continual learning? Due to the computational intensity of DFL, cumulative retraining or replaying past tasks leads to prohibitive computational costs under bounded model and memory capacity [24]. ii) How to ensure the computationally feasibility? The mixed-integer coupling of logistics decisions often yields discontinuous gradients in back-propagation of DFL and destabilizes end-to-end training [25].
@@ 6 Conclusion
This work addresses the task-generalization challenge of decision-focused forecast models in tasks streams of PLS scheduling by introducing a DFCL paradigm. The proposed DFCL removes the need to retrain a bespoke model for every new task by applying parameter regularization to mitigate forgetting. Across sequences of tasks with varying operational parameters, DFCL consistently improves decision quality and generalization over naïve Adam-based continual learning and markedly reduces training cost relative to task-wise DFL and multi-task retraining. This work provides a practical path for scalable deployment of forecasting models in dynamic seaport operations. Future research could explore the theoretical underpinnings of task relatedness in DFCL to deepen understanding and broaden applicability to other predict-and-optimize domains.

### v2

# Predict-then-Optimize for Seaport Power–Logistics Scheduling: Generalization across Varying Tasks Stream

@@ Abstract
Power–logistics scheduling in modern seaports typically follow a predict-then-optimize pipeline. To enhance the decision quality of forecasts, decision-focused learning has been proposed, which aligns the training of forecasting models with downstream decision outcomes. However, this end-to-end design inherently restricts the value of forecasting models to only a specific task structure, and thus generalize poorly to evolving tasks induced by varying seaport vessel arrivals. We address this gap with a decision-focused continual learning framework that adapts online to a stream of scheduling tasks. Specifically, we introduce Fisher information based regularization to enhance cross-task generalization by preserving parameters critical to prior tasks. A differentiable convex surrogate is also developed to stabilize gradient backpropagation. The proposed approach enables learning a decision-aligned forecasting model across a varying tasks stream with a sustainable long-term computational burden. Experiments calibrated to the Jurong Port demonstrate superior decision performance and generalization over existing methods with reduced computational cost.
@@
[2]organization=School of Electrical and Electronic Engineering, Nanyang Technological University, country=Singapore
@@ 1 Introduction
To address this misalignment, decision-focused learning (DFL) instead evaluates the decision quality of forecasts through their impact on the downstream optimization tasks [8]. As depicted in Fig. 1(b), DFL integrates forecasting and optimization in an end-to-end pipeline and trains the forecasting model directly by the decision-related loss. Recent work further refines DFL for power system applications by deriving task-aware losses and differentiable solvers that facilitate stable end-to-end training. These studies have established an effective learning paradigm for enhancing decision-making benefits in various power system contexts, including unit commitment [9, 10], economic dispatch [12, 13, 14], cost-oriented renewable-energy forecasting and trading [15, 16, 17], energy-storage arbitrage [18, 19], and multi-energy scheduling [20].
However, this end-to-end learning paradigm inherently confines the forecasting model’s decision quality to a fixed and predetermined optimization task. In contrast, PLS scheduling tasks are continuously evolving due to varying vessel arrivals, since logistics decisions must be made conditional on vessel arrivals as exogenous parameters [7]. Each new arrival configuration introduces a distinct task instance characterized by different numbers of vessels, cargo profiles, and time-coupled constraints. Although these tasks share similar high-level formulations, they differ in the dimensions of decision variables and the geometry of feasible regions. As a result, the ground-truth objectives and loss landscapes of DFL is correspondingly varying across different configurations. This constitutes the fundamental distinction between PLS operation and conventional DFL settings in power systems.
@@
Online continual learning thus provides a computationally sustainable solution for long-term adaptation, yet two challenges still remain: i) How to retain the decision quality of forecasts on past tasks? Due to the computational intensity of DFL, cumulative retraining or replaying past tasks leads to prohibitive computational costs under bounded model and memory capacity [24]. ii) How to ensure the computationally feasibility? The mixed-integer coupling of logistics decisions often yields discontinuous gradients in back-propagation of DFL and destabilizes end-to-end training [25].
@@ 6 Conclusion
This work addresses the task-generalization challenge of decision-focused forecasting in the sequential task stream of PLS scheduling. The proposed DFCL framework eliminates the need to retrain a bespoke model for each new task. Across sequences of tasks with varying operational parameters, DFCL consistently enhances decision quality and cross-task generalization compared with naïve Adam-based continual learning, while markedly reducing the computational cost relative to task-wise DFL and multi-task retraining. Rather than providing a one-shot solution that addresses a fixed set of tasks, the proposed framework continuously adapts to evolving task conditions. This offers a practical and efficient approach for seaports to maintain high-quality scheduling decisions amid changing operational contexts. Future research could explore the theoretical underpinnings of task relatedness in DFCL to deepen understanding and broaden applicability to other predict-and-optimize domains.